OpenStudy (anonymous):

how would you do 5 2 - + _ 6 3

OpenStudy (anonymous):

you have to find a common denominator

OpenStudy (anonymous):

in this case it would be 6

OpenStudy (anonymous):

Thank You @Austin6i6

OpenStudy (anonymous):

Example \[\frac{3}{6} + \frac{1}{3} \] \[\frac{3}{6} + \frac{1}{3}*\frac{2}{2}\] \[\frac{3}{6}+\frac{2}{6}\]